# Regimes of Steady-State Turbulence in a Quantum Fluid

### v1

@@ V.1 Discussion
Weak wave turbulence.— For weaker forcing, $0.4\mu\lesssim U_{F}\lesssim 1.2\mu$, there is a $\propto k^{-3.5}$ power law spectrum for compressible kinetic energy, quantum pressure, and the total spectrum, and a distinct absence of extended vortex lines in the bulk superfluid. In this weak-wave turbulence regime $E^{i}_{\text{kin}}/E^{c}_{\text{kin}}\ll 1$. Even when small vortex rings appear in the steady state as $U_{F}$ approaches $\mu$, the $k^{-3.5}$ scaling is robust and spans a decade of $k$-space. The forcing amplitude reported in Ref. [5], $U_{F}=0.8\mu$, is in the center of the $U_{F}$ range where we identify the scaling both instantaneously [Fig. 4] and in time averages [Fig. 5]. In addition, we observe a clear forcing peak near the system scale $L_{z}$ in the total wave-action spectrum at early times. The wave-action spectra decomposition, combined with absence of bulk vortices implicates a direct Bogoliubov quasiparticle cascade for the $k^{-3.5}$ power law. Indeed, Bogoliubov phonons involve a quantum pressure component associated with small scale density fluctuations when $k\xi\gtrsim 1$. In Fig. 5(b), we see $n^{c}(k)$ and $n^{q}(c)$ are closely matched within this $k$ range, and decouple when $k\xi\lesssim 1$.
The log-corrected spectrum analysis of WWT direct cascade is strictly applicable for very weak forcing in the absence of a condensate [19], and is thus not directly applicable to our system that instead involves a large condensate with dominant three wave interactions. However, the three wave kinetic equation also supports a $k^{-3}$ scaling [21]. Applying the log-correction to our wave-action spectra, we find good agreement if we identify the forcing scale for phonons as $k_{F}=1/\xi$, instead of the natural forcing wavenumber of the potential $2\pi/L_{z}$ as detailed in Appendix C. Indeed, the analysis of the forcing scale by Martirosyan et al[23], reached a similar conclusion  [^6] Their result translates to $k_{F}=\sqrt{2}/\xi$ in our units. However, for this estimate of the forcing scale we see a narrower flat region in the log-corrected spectrum. For further discussion see Appendix C..
@@
Strong vortex turbulence.— For $2.5\mu\lesssim U_{F}$, the condensate bulk becomes saturated with vortex lines and the compressible spectrum approximates a $k^{-7/3}$ power-law over more than a decade of $k$; this is consistent with the predicted WWT scaling for an inverse particle-cascade solution to the four-wave kinetic equation. In Ref. [19], this was observed in the total spectrum for weak forcing in the absence of BEC, while here we see it in the compressible spectrum only, for strong forcing. As the power law extends far below $k_{\xi}=1/\xi$, this would be consistent with WWT if random vortices effectively destroy the condensate, and indeed the phase coherence length drops to the healing length at $U_{F}\gtrsim 2\mu$ [Fig. 6]. As our model only contains damping for high wavenumbers, a steady inverse-cascade of particles can only exist in $n^{c}(k)$ if particles are transported to high $k$ by the other components of $n(k)$. Physical consistency thus requires a compressible particle source near the dissipation scale $k\sim k_{D}$ to generate an inverse cascade; one candidate mechanism is the steady annihilation of vortices, creating bursts of sound at small scales. In Figure 3(d),(e) the compressible energy driving by the forcing appears to become ineffective for $U_{F}\gtrsim 2\mu$, supporting the interpretation that compressible energy is instead injected at small scales via vortex annihilation.
@@ V.2 Conclusions
Motivated by recent experiments in forced quantum turbulence [5, 15, 9], we have simulated a Gross-Pitaevskii model for a wide range of forcing energies, generating and analysing steady turbulence. We find evedence for regimes of weak-wave and strong-vortex turbulence separated by a mixed regime that lacks clear power-law behavior. Distinct regimes of power-law behavior are identified by application of high resolution spectral analysis [24], and further studied through vortex distributions and two-point correlations. Our results are largely consistent with other recent works [23], while also enabling further decomposition of the wave-action spectrum. For weak forcing, the bulk superfluid is approximately vortex free, and exhibits a power-law of $\sim k^{-3.5}$, consistent with weak-wave turbulence predictions for a direct cascade [19]. The compressible and quantum pressure components also show the same power-law scaling and are in close agreement. Log-correction analysis implicates a forcing wavenumber $k_{F}\sim\xi^{-1}$ for the cascade. In this regime the superfluid order parameter remains coherent over the system scale, with coherent only weakly degraded by small vortex rings. In contrast, for strong forcing the steady state develops a dense tangle of extended vortices that limites phase coherence to the healing length scale. Power-law scaling $\sim k^{-7/3}$ is observed in the compressible component only, consistent with the weak-wave turbulence prediction of an inverse cascade in non-condensate four-wave kinetics [19].
@@ Acknowledgements
We thank Gevorg Martirosyan and Tim Copland for stimulating discussions. TF would like to thank the University of Otago for financial support.
@@ B.1 Calculation of spectral densities using Plancherel’s theorem
Figure 9: Comparison of velocity power spectra and energy spectral densities. (a) Compressible components. The two spectra are entirely dissimilar, except for a small central peak. (b) Incompressible components. The two distributions show strong agreement in the inertial range, but are distinct in the ultraviolet and infrared. This suggests that the turbulence is semiclassical, corroborating the random phase approximation
@@ Appendix C Comparison with Logarithmic-correction model
Another explanation for the $k^{-3.5}$ power law, is that the $k\propto k^{-3}$ stationary solution is marginally non-local, requiring a log-correction of the form
@@
This logarithmic-correction emerges from the introduction of an infrared cutoff at the forcing wavenumber, $k_{F}$. In Ref. [19], $k_{F}$ is a defined parameter. With our method of energy injection this is not the case, and $k_{F}$ is slightly more ambiguous. We excite the condensate along the $z$-direction, so one might expect $k_{F}=k_{L_{z}}$. However, in practice we find the forcing peak appears at slightly higher $k$-values than $k_{L_{z}}$ [Fig. 4], implying the forcing length scale is slightly smaller than $L_{z}$. Regardless, we test the models fit using a wide range of $k_{F}$ values.
@@
Applying the logarithmic-correction to spectra from the $k^{-3.5}$ regime, we generally find poor agreement using $k_{F}$ values near the actual forcing lengthscale. However, for $k_{F}=1/\xi$, we find good agreement in the total momentum distribution and slightly stronger agreement in the compressible & quantum pressure components [Fig. 10]. This $k_{F}$ value is clearly distinct from the actual forcing wavenumber, supporting the argument presented in Ref. [23], that the physical energy-injection lengthscale is distinct from the effective injection scale for the four-wave cascade. The strength of this agreement, however, is limited by the size of the inertial range, which to create a more compelling case could be increased with a greater dissipation lengthscale and a finer numerical grid to ensure the model remains stable.

### v2

@@ V.1 Discussion
Weak wave turbulence.— For weaker forcing, $0.4\mu\lesssim U_{F}\lesssim 1.2\mu$, there is a $\propto k^{-3.5}$ power law spectrum for compressible kinetic energy, quantum pressure, and the total spectrum, and a distinct absence of extended vortex lines in the bulk superfluid. In this weak-wave turbulence regime $E^{i}_{\text{kin}}/E^{c}_{\text{kin}}\ll 1$. Even when small vortex rings appear in the steady state as $U_{F}$ approaches $\mu$, the $k^{-3.5}$ scaling is robust and spans a decade of $k$-space. The forcing amplitude reported in Ref. [5], $U_{F}=0.8\mu$, is in the center of the $U_{F}$ range where we identify the scaling both instantaneously [Fig. 4] and in time averages [Fig. 5]. In addition, we observe a clear forcing peak near the system scale $L_{z}$ in the total wave-action spectrum at early times [Fig. 4]. The wave-action spectra decomposition, combined with absence of bulk vortices implicates a direct Bogoliubov quasiparticle cascade for the $k^{-3.5}$ power law. Indeed, Bogoliubov phonons contain a quantum pressure component associated with small length-scale density fluctuations when $k\xi\gtrsim 1$. In Fig. 5(a), we see $n^{c}(k)$ and $n^{q}(c)$ are closely matched within this $k$ range, and decouple when $k\xi\lesssim 1$. The role of quantum pressure may be seen by linearizing the GPE around a homogeneous condensate and using the Madelung decomposition [25]. For Hamiltonian GPE evolution, the superfluid Euler equation together with continuity can be used to derive the equation of motion for linearized density fluctuations $\delta n\equiv n-n_{0}$. They obey the generalized wave equation
$\frac{\partial^{2}\delta n}{\partial t^{2}}=\frac{gn_{0}}{m}\nabla^{2}\delta n-\left(\frac{\hbar}{2m}\right)^{2}\nabla^{4}\delta n.$ (31)
The first term corresponds to the hydrodynamic regime, while the second term generates the free-particle behavior at high $k$ and stems from the quantum pressure term in the GPE. Assuming plane wave solutions for the density $\delta\rho\sim\Re\{\delta\rho_{0}\exp{[i(kr-\omega t)]}\}$, we recover the Bogoliubov dispersion relation
$\hbar\omega=\sqrt{\frac{\hbar^{2}k^{2}}{2m}\left(2gn_{0}+\frac{\hbar^{2}k^{2}}{2m}\right)}.$ (32)
A phonon cascade at high $k$ will necessarily involve significant quantum pressure, due to its role in short wavelength quasiparticle excitations.
The log-corrected spectrum analysis of WWT direct cascade is derived for very weak forcing in the absence of a condensate [19], and is thus not directly applicable to our system that involves a large condensate and hence three-wave interactions. However, the three-wave kinetic equation also supports a $k^{-3}$ power law [21]. Furthermore, we can expect four-wave interactions to dominate when either (i) the condensate is highly depleted due to strong forcing, or (ii) the interaction is centered around high $k$ values where there is no condensate. Applying the log-corrected $k^{-3}$ form to our wave-action spectra, we find good agreement provided we identify the forcing wavenumber for phonons as $k_{F}=1/\xi$, instead of the natural forcing wavenumber of the potential $2\pi/L_{z}$ as detailed in Appendix C. Our findings are consistent with the analysis of Martirosyan et al[23], who reached a similar conclusion  [^7] Their result translates to $k_{F}=\sqrt{2}/\xi$ in our units. However, for this estimate of the forcing scale we see a narrower flat region in the log-corrected spectrum. For further discussion see Appendix C..
@@
Strong vortex turbulence.— For $2.5\mu\lesssim U_{F}$, the condensate bulk becomes saturated with vortex lines and the compressible spectrum approximates a $k^{-7/3}$ power-law over more than a decade of $k$; this is consistent with the predicted WWT scaling for an inverse particle-cascade solution to the four-wave kinetic equation. In Ref. [19], this was observed in the total spectrum for weak forcing in the absence of BEC, while here we see it in the compressible spectrum only, for strong forcing. As the power law extends far below $k_{\xi}=1/\xi$, this would be consistent with WWT if random vortices effectively destroy the condensate, and indeed the phase coherence length drops to the healing length at $U_{F}\gtrsim 2\mu$ [Fig. 7]. As our model only contains damping for high wavenumbers, a steady inverse-cascade of particles can only exist in $n^{c}(k)$ if particles are transported to high $k$ by the other components of $n(k)$. Physical consistency thus requires a compressible particle source near the dissipation scale $k\sim k_{D}$ to generate an inverse cascade; one candidate mechanism is the steady annihilation of vortices, creating bursts of sound at small scales. In Figure 3(d),(e) the compressible energy driving by the forcing appears to become ineffective for $U_{F}\gtrsim 2\mu$, there is also negligable variation in spectrum amplitude with forcing [Fig. 6], supporting the interpretation that compressible energy is instead injected at small scales via vortex annihilation.
@@ V.2 Conclusions
Motivated by recent experiments in forced quantum turbulence [5, 15, 9], we have simulated a Gross-Pitaevskii model for a wide range of forcing energies, generating and analyzing steady turbulence. We find evidence for regimes of weak-wave and strong-vortex turbulence separated by a mixed regime that lacks clear power-law behavior. Distinct regimes of power-law behavior are identified by application of high resolution spectral analysis [24], and further studied through vortex distributions and two-point correlations. Our results are largely consistent with other recent works [23], while also enabling further decomposition of the wave-action spectrum. For weak forcing, the bulk superfluid is approximately vortex free, and exhibits a power-law of $\sim k^{-3.5}$, consistent with weak-wave turbulence predictions for a direct cascade [19]. The compressible and quantum pressure components also show the same power-law scaling and are in close agreement. Log-correction analysis implicates a forcing wavenumber $k_{F}\sim\xi^{-1}$ for the cascade. In this regime the superfluid order parameter remains coherent over the system scale, with coherent only weakly degraded by small vortex rings. In contrast, for strong forcing the steady state develops a dense tangle of extended vortices that limits phase coherence to the healing length scale. Power-law scaling $\sim k^{-7/3}$ is observed in the compressible component only, consistent with the weak-wave turbulence prediction of an inverse cascade in non-condensate four-wave kinetics [19].
@@ Acknowledgements
We thank Gevorg Martirosyan, Nir Navon, and Tim Copland for stimulating discussions. TF would like to thank the University of Otago for financial support.
@@ B.1 Calculation of spectral densities using Plancherel’s theorem
Figure 10: Comparison of velocity power spectra and energy spectral densities. (a) Compressible components. The two spectra are entirely dissimilar, except for a small central peak. (b) Quantum Pressure, where the two spectra are again distinct. (c) Incompressible components. The two distributions show strong agreement in the inertial range, but are distinct in the ultraviolet and infrared. This suggests that the turbulence is semiclassical, corroborating the random phase approximation
@@ Appendix C Comparison with Logarithmic-correction model
Another explanation for the $k^{-3.5}$ power law is that the $k\propto k^{-3}$ stationary solution is marginally non-local [19], requiring a log-correction of the form
@@
This logarithmic-correction emerges from the introduction of an infrared cutoff at the forcing wavenumber, $k_{F}$. In Ref. [19], $k_{F}$ is a defined parameter. With our method of energy injection this is not the case, and $k_{F}$ is slightly more ambiguous. We excite the condensate along the $z$-direction, so one might expect $k_{F}=k_{L_{z}}$. However, in practice we find the forcing peak appears at slightly higher $k$-values than $k_{L_{z}}$ [Fig. 4], implying the forcing length scale is slightly smaller than $L_{z}$. Regardless, we test the model fit using a wide range of $k_{F}$ values.
Applying the logarithmic-correction to spectra from the $k^{-3.5}$ regime, we generally find poor agreement using $k_{F}$ values near the actual forcing lengthscale. However, for $k_{F}=1/\xi$, corresponding to a length scale of $2\pi\xi$, we find good agreement in the total momentum distribution and slightly stronger agreement in the compressible and quantum pressure components [Fig. 12]. This $k_{F}$ value indicates a forcing scale distinct from both the obvious forcing scale $L_{z}$ from the drive, and the healing length. In Ref. [23] similar analysis showed that for 4-wave interactions the physical forcing scale due to the drive is distinct from the effective injection scale for weak wave turbulence.